What does a graph of 16y^2-9x^2=144 look like? For a parabola, state the vertex, the focus, and the equation of the directrix. For a circle, state the center and radius. For an ellipse, state center, vertices, co-vertices, and foci. For a hyperbola, state the center, vertices, and foci.
Answer by lwsshak3(11628)

What does a graph of 16y^2-9x^2=144 look like? For a parabola, state the vertex, the focus, and the equation of the directrix. For a circle, state the center and radius. For an ellipse, state center, vertices, co-vertices, and foci. For a hyperbola, state the center, vertices, and foci.
***
16y^2-9x^2=144
divide both sides by 144
y^2/9-x^2/16=1
This is an equation of a hyperbola with vertical transverse axis
center: (0,0)
a^2=9
a=3
b^2=16
b=4
c^2=a^2+b^2=9+16+25
c=5
vertices:(0,0±a)=(0,0±3)=(0,3) and (0-3)
foci:(0,0±c)=(0,0±5)=(0,5) and (0-5)